Suppose the total product (Q) is given by the production function; [q = 4KL-0.5]

Where K= capital = 12 and L = unit of labour -36.

Calculate the average product of labour


1
Expert's answer
2021-03-23T08:29:26-0400

We can find the average product of labour as follows:


"AP_L=\dfrac{q}{L},""AP_L=\dfrac{4KL^{-0.5}}{L}=\dfrac{4K}{L^{1.5}}."

Substituting the numbers, we get:


"AP_L=\dfrac{4K}{L^{1.5}}=\dfrac{4\cdot12}{36^{1.5}}=0.22"